# Good News for Polynomial Root-finding

### v35

@@ 1.8.1 Our novelties (a short list)
Some of our techniques for black box polynomial root-finding, e.g., randomization for root-finding, numerical stabilization of root-squaring, acceleration of black box subdivision root-finding by means of compression of a disc without losing any zero of $p$, acceleration of bisection by means of applying it to the exponent range, approximation of root radii and near real roots, and even our simple but novel recipes for detection of output errors of polynomial root-finders can be of independent interest.
1.8.2 Related works
Unlike deterministic solution in [92, 99] we use Las Vegas randomization but propose algorithms that work for a black box polynomial, are simpler to implement, and in the case of a polynomial $p$ given with its coefficients support the same near optimal complexity bound of [92, 99].
@@ 1.8.2 Related works
The pioneering paper [73] proposed a novel algorithm for approximation of an absolutely largest zero of a real-rooted black box polynomial, estimated the computational precision of that algorithm, and extended it to approximation of an absolutely largest eigenvalue of a symmetric matrix. We adopt their model of computation, their extension to approximation of matrix eigenvalues, and some of their patterns for estimating the computational precision of root-finders, but otherwise use distinct approach and techniques.
Namely, we accelerate the classical subdivision root-finders, based on e/i tests. All known e/i test operate with the coefficients of $p$: they involve higher order derivatives $p^{(i)}(x)$, $i=1,2,\dots,d$ [50, 124], Newton’s identities and fast root-squaring algorithms based on convolution [97], or Pellet’s theorem [22, 23]. Neither of these approaches works for black box polynomials.
Our soft e/i tests involve the classical root-squaring techniques of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837, but we stabilize them numerically avoiding notorious problems that severely limit all previous applications of maps (1.4) of shift and scaling and (1.5) of DLG iterations.
Renegar in [124] accelerated subdivision process by a factor of $b/\log(b)$ by incorporating Newton’s root-finding iterations; the papers [97, 22, 23] combined his approach with alternative functional root-finding iterations. We apply a distinct technique of compression of a disc instead.
@@
Estimating the bit operation complexity of our root-finders we use [130, Eqn. (9.4)] and [78, Thm. 2] for multipoint polynomial evaluation. [^16] The ingenious algorithm of Moroz supporting his [78, Thm. 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$. To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$. At this point, application of [65, Alg. 5.3 and Thm. 3.9] yielded new record fast algorithm for multipoint polynomial approximation.
In Sec. 11.4 we accelerate e/i tests by a factor of $m$ under a random root model and accordingly accelerate the solution of Problems 1-3 and $*{}^{*}$$*{}^{*}$ by such a factor. Moroz in [78] and Imbach and Moroz in [56] accelerated their solution under an upper bound on condition numbers of all zeros of $p(x)$ instead. In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [13, 28, 21], this is a little stronger restriction than bounding cluster sizes of Def. B.1, under which [92, 99] proved such a complexity bound.
Without using the algorithm of [78, Thm. 2], our current algorithms still support such a near optimal bit operation cost bound for Problem 1 with no restriction on an input polynomial $p$ in the case where $m^{2}=\tilde{O}(d)$, $1\leq m\leq d$ (see [108, Remark 20.1]) as well as under the Random Root Model of Sec. 1.5.
@@ 1.9 Implementation: State of the Art
The algorithm of [130] was implemented by Xavier Gourdon in [44] for the Magma and PARI/GP computer algebra systems, while the near optimal but more involved algorithm of [92, 99] has never been implemented.
@@
Our current algorithms are much simpler for implementation than that of [92, 99] and, if properly implemented, can become user’s choice root-finders. Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder of [61] has enabled its noticeable acceleration (see [57]). The more advanced publicly available implementation in [60] “relies on novel ideas and techniques" from [108] (see [60, Sec. 1.2]). According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [61] or any other known variant of complex subdivision root-finder, has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^17] The complexity of a subdivision root-finder in a region is proportional to the number of roots in it, while MPSolve is about as fast and slow for all roots as for their fixed subset. MPSolve implements Ehrlich-Aberth’s root-finding iterations, which empirically converge to all $d$ roots very fast right from the start but only under empirical support and only under initialization that operates with the coefficients of $p$ [125]. and tended to supersede MPSolve significantly as the degree $d$ of $p$ grew large. Such a performance could only be a dream for the designers of subdivision root-finders before 2022, but there is a large room for further improvement because [60] still inherits some earlier algorithms from [22, 61], which are slower by order of magnitude than our current ones (see some specific directions next).
1.10 Tentative directions for further progress
Next we comment on some specific directions for significant (and perhaps dramatic) acceleration of the root-finders of [57, 60], based partly on our present work and partly on its expected further advance. Since our algorithms are near optimal, we only seek relaxation of the assumptions of the Random Root Model and empirical acceleration.
(A) Acceleration of e/i tests. They are the most used ingredient of our subdivision root-finders as well as those of [57, 60].
(i) In Part III we by-pass approximation of the Cauchy integral of (1.9) and only evaluate the polynomial $p$ at or near the unit circle $C(0,1)$ (see Sec. 11.4).
(ii) We achieve additional acceleration under the Random Root Model because we only need to perform root-lifting and then to evaluate $p$ at a small number of random points $x$. The cost of root-lifting is dominated in our formal study but can substantially slow down implementation. We conjecture that empirically the algorithm should work without lifting or with a small number of lifting steps.
(iii) E/i tests by means of approximation of the distance from a complex point to the roots seems to be competitive or maybe even superior empirically (see Remark 1.2).
(iv) We can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [141]), efficiently implemented in Boost C++ libraries. This promises no significant acceleration for the worst case inputs but may be useful empirically.
(B) Acceleration of compression of a disc.
We accelerate the compression in [57, 60] at two stages:
(i) The algorithms of Secs. 20.2 and 20.3 are much faster than those of [57, 60], based on Cauchy sum, for computation of a point that lie in or near the convex hull of the root set in a fixed disc.
(ii) The algorithm of Sec. 12.1 runs significantly faster than those in [57, 60] for approximation of the smallest root radius from a given point on the complex plane, and those based on the DLG iterations (see Remark 1.2) are potentially competitive. Notice that we specified and analyzed extension from approximation of the distance to roots to approximation of the distance to the $m$ closest root in a disc containing precisely $m$ roots (see Sec. 12).
(C) Acceleration of real and near real root-finding. Our novel algorithm of Sec. 9 can be readily implemented and seems to be the first that efficiently solves this important problem.
(D) Acceleration of approximation of matrix eigenvalues. We adopt extension of root-finding to proposed in [73] because it implies record low upper estimates in the case of matrices and matrix polynomials of very large size, but this extension relies on the computation of the coefficients of the characteristic polynomial of an input matrix or matrix polynomial, which is prone to numerical stability problems and avoided in practice of numerical matrix computations. In Remark 7.1 of Sec. 7 we point out an alternative extension, which avoids such a notorious stage and is dramatically faster where the input matrix or matrix polynomial is data sparse.
@@ 1.11 Organization of the rest of the paper
In Parts II and III we devise and analyze our two alternative e/i tests. Deterministic e/i test of Part II is slower than that of Part III, but the root-finders of both parts use Las Vegas randomization to bound computational precision and bit-operation cost. These bounds are roughly the same in both parts, except thatin Part III we yield acceleration by a factor of $d$ under the Random Root Model. All the bit-operation complexity bounds of Parts II and III are is inferior by a factor of $b$ to that of Thm. 1.3.
We complete the proof of Thm. 1.3 in Part IV by means of our novel acceleration of subdivision iterations based on compression of a disc without losing its roots.
In Sec. A of the Appendix we comment on applications and implementation of root-finding. In its Secs. B and C we outline the fundamental results of Schönhage from [130] on polynomial factorization, its links to root-finding, and root-isolation and some relevant bit-complexity estimates. In Sec. D we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and strengthen them with incorporation of Fast Multipole Algorithm (FMM).
@@ 1.12 Independence of Part III from Part II
In Parts II an III we solve Problems 0, 1, $*{}^{*}$, and $*{}^{*}$ in two different ways. In Part II we rely on approximation of a Cauchy integral associated with root-counting in a disc on the complex plane. In Part III we rely on lifting the zeros of $p$ to a fixed power $h=O(\log(d))$. In addition, in Part III, unlike Part II, we devise a faster root-finder under a random root model. We make our presentation in Part III independent of Part II; in particular we include again some definitions, figures, and theorems.
@@
[…]
@@ 2 Background: basic definitions
  • $\Delta(\mathbb{R})$, $X(\mathbb{R})$, $CH(\mathbb{R})$, and $\#(\mathbb{R})$ denote the diameter, the root set, the convex hull, and the index (root set’s cardinality) of a region $\mathbb{R}$ on the complex plane, respectively.
  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is said to be $\theta$-isolated for $\theta>1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho/\theta))$, or $X(S(c,\rho))=X(S(c,\theta\rho))$, respectively.
@@
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex center $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. $r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
[…]
Observation 2.1.
@@ 3.2 Cauchy sums in the unit disc as weighted power sums of the roots
Cor. 3.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
@@
We can apply it because $|y|=|z|$ while $|z|<1$ by assumption.
@@
Sum equations (3.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 1.13 and (3.2), and obtain
@@
For the zeros $z_{j}$ of $p(x)$ and all $h$, the Cauchy sums $s_{h,q}$ of (3.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}~{}{\rm unless}~{}z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$

### v36

@@ 5.1 Our novelties (a short list)
Some of our techniques such as randomization, numerical stabilization of root-squaring, compression of a disc without losing roots, bisection of logarithms, approximation of root radii and near-real roots, and even our simple error detection recipe can be of independent interest.
5.2 Related works
@@
The pioneering paper [76] proposed a novel algorithm for approximation of an absolutely largest zero of a real-rooted black box polynomial, estimated the computational precision of that algorithm, and extended it to approximation of an absolutely largest eigenvalue of a symmetric matrix. From that paper we adopt the model of computation, extension to approximation of matrix eigenvalues, and some patterns for estimation of the computational precision of root-finders, but otherwise use distinct approach and techniques.
Namely, instead of Newton’s iterations in [76], we accelerate the classical subdivision root-finders, based on e/i tests. Its closest variations in [128, 101, 23, 24] still use inferior e/i tests, for which we proposed two fast novel approaches – based on approximation of Cauchy integrals and on root-lifting; these tests run much faster than the tests in the cited papers and, unlike the latter tests, can be applied to a black box polynomial. The papers [128, 101, 23, 24] apply functional iterations (such as Newton’s and Schröders) to accelerate convergence of subdivision process but we apply our novel techniques of disc compression instead.
@@
Estimating the bit operation complexity of our root-finders we use [134, Eqn. (9.4)] and [81, Thm. 2] for multipoint polynomial evaluation. [^8] The ingenious algorithm of Moroz supporting his [81, Thm. 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$. To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$. At this point, application of [68, Alg. 5.3 and Thm. 3.9] yields new record fast algorithm for multipoint approximation of a polynomial given with its coefficients. The root-finders of Moroz in [81] and Imbach and Moroz in [59] use $\tilde{O}(bd)$ bit operations for polynomials given with their coefficients and having only well-conditioned zeros. In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [14, 30, 22], the upper bound on the condition numbers of all zeros of $p(x)$ in [81, 59] is a little stronger restriction than the bound on the cluster sizes of Def. 24.1, under which [96, 103] proved the same complexity bound. Unlike our present root-finders, the root-finders of [81, 59] are not efficient for the worst case inputs and extensively use the coefficients.
@@ 5.3 Implementation: State of the Art
The near-optimal but quite involved algorithm of [96, 103] has never been implemented, while the slower algorithm of [134] has been implemented by Xavier Gourdon in [46] for the Magma and PARI/GP computer algebra systems. By now it is superseded by implementation of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (in Sec. 26.4 we list some competitive implementations of Newton’s iterations).
Current user’s choice for real root-finding is the algorithm of [70], although it has already been superseded in [62] according to extensive tests for standard test polynomials.
@@
We present algorithms that are rather simple for implementation and promise to compete for user’s choice. Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder [64] has enabled its noticeable acceleration (see [60]). The more advanced publicly available implementation in [63] “relies on novel ideas and techniques" from [112] (see [63, Sec. 1.2]). According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [64], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^9] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [96, 103] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots. Thus already implementation in [64] of subdivision iterations, slightly outperformed MPSolve for root-finding in a domain $\mathbb{D}$ containing a small number of roots although is by far inferior for the approximation of all $d$ roots. and tended to supersede MPSolve more and more significantly as the degree $d$ of $p$ grew large.
There is still a large room for further improvement because the root-finder of [63] inherits some earlier sub-algorithms from [23, 64], which are slower by order of magnitude than our current ones.
Next we comment on some promising directions to further improvement of [63], but also notice that adjustment of our near-optimal subdivision root-finder in Remark 2.1 to approximation of real and near-real roots and consequently of matrix eigenvalues should be at least competitive with the current user’s choice [70] for real root-finding as well as with its significant improvement in [62]; moreover, it more efficiently counters the impact of the input and rounding errors.
5.4 Further progress and promise
(A) Acceleration of e/i tests. E/i tests are basic ingredient of subdivision root-finders. [60, 63] used the Cauchy-based e/i tests of our Part II and their heuristic acceleration. It is still a challenge to compare their empirical performance with that of the formally supported e/i tests of our Part III and Sec. 23.4, based on root-lifting. Furthermore, we can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [146]), efficiently implemented in Boost C++ libraries.
(B) Acceleration of the compression of a disc. We assume that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the external $d-m$ roots, and we compress that disc towards the minimal disc containing the same $m$ roots. We reduce this task essentially to (i) computing a point in or near the convex hull of the root set in $D(0,1)$ and (ii) approximation of the $m$th smallest root radius from a given point in $D(0,1)$. For task (i) we dramatically accelerate our algorithm of [60, 63] because our algorithms of Secs. 23.2 and 23.3 for task (i), using Schröder’s iterations, deterministic, and formally supported, run dramatically faster than those of [60, 63], based on approximation of the average of all roots in $D(0,1)$. For task (ii) we can apply our Las Vegas randomized algorithms of Secs. 8 and 9 combined or the alternative of [60, 63], based on the algorithm of our Sec. 15.1, which reduces this task to performing a constant number of e/i tests. Both approaches support the same asymptotic cost bound up to polylog factors, but fro $m\gg\log^{2}(d)$ the second approach is faster by a factor of $m/\log^{2}(d)$ where it uses fast heuristic e/i tests. Such an improvement, however, may be lost in practice because of the impact of polylog factors. Implementation of compression in [60, 63] is slowed down with verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our Las Vegas root-finders.
(C) Fast approximation of the eigenvalues of a matrix. [20] explored reduction of approximation of the eigenvalues of a matrix to root-finding by means of Ehrlich’s iterations, reduced to recursive computations of NIR$(x)$, and showed benefits of this approach. We can readily extend this reduction and its benefits by using subdivision iterations instead of Ehrlich’s. Furthermore, win the case of subdivision we obtain additional benefits, e.g., our algorithms are much faster for approximation of a small number of eigenvalues that lie in a fixed domain and unlike Ehrlich’s iterations isolate eigenvalues, rather than just approximate them.
(D) Acceleration of real and near real root-finding. In Remark 2.1 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors and can be readily extended to approximation of the eigenvalues of a matrix.
(E) Acceleration by using fast multipoint polynomial evaluation. Mere incorporation of fast algorithms of [81, 59] should greatly accelerate the performance of [60, 63] in the case of root-finding for polynomials given with their coefficients.
@@ 5.5 Organization of the rest of the paper
In Parts II and III we devise and analyze our two alternative e/i tests. Deterministic e/i test of Part III is as fast as the randomized one of Part II. We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost, and in Part III we yield acceleration by a factor of $d$ under the Random Root Model. The bit-operation complexity bounds of Parts II and III are inferior by a factor of $b$ to that of [96, 97, 103], but we match it in Part IV by means of our acceleration of the algorithms of both Parts II and III based on compression of a disc without losing its roots. In Part V we cover some important root-finding subjects that involve no substitution iterations, namely, factorization, root isolation, and functional iterations. In Secs. 24 and 25 we outline the fundamental results of Schönhage from [134] on polynomial factorization, its links to root-finding and root-isolation, and some relevant bit-complexity estimates, including lower bounds. In Sec. 26 we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and their acceleration by means of incorporation of Fast Multipole Algorithm (FMM) and fast multipoint polynomial evaluation as well as their initialization both where the coefficients of $p$ are unknown or given to us.
@@ 5.6 Independence of Part III from Part II
In Parts II an III we solve Problems 0, 1, 0, and 1 in two different ways. In Part II we rely on approximation of a Cauchy integral associated with root-counting in a disc on the complex plane. In Part III we rely on lifting the roots to a fixed power $h=O(\log(d))$. In addition, in Part III, unlike Part II, we devise a faster root-finder under a random root model. We make our presentation in Part III independent of Part II; in particular we include again some definitions, figures, and theorems.
@@
Organization of the presentation in Part II: We devote the next section to some basic definitions. In Sec. 7 we study root-counting in a disc on the complex plane by means of approximation of a Cauchy integral over its boundary circle by finite sums. [^10] We do not use the results of Sec. 7.5, but they can be of independent interest. In Sec. 8 we apply this study to devise and analyze root-counters and e/i tests in a disc on the complex plane. In Sec. 9 we extend our e/i tests to estimating root radii, that is, the distances from a fixed complex point to the roots. In Secs. 10 we estimate precision of computing of our root-finders. We devote short Sec. 11 to their extension to approximation of the eigenvalues of a matrix or a matrix polynomial. In Sec. 12 we estimate the bit operation complexity of our root-finders assuming that $p(x)$ is a general polynomial of (1.1) represented with its coefficients.
@@ 6 Background: basic definitions
  • $\Delta(\mathbb{S})$, $X(\mathbb{S})$, $CH(\mathbb{S})$, and #$\#(\mathbb{S})$ denote the diameter, the root set, the convex hull, and the index (root set’s cardinality) of a set $\mathbb{S}$ in the complex plane, respectively.
  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is said to be $\theta$-isolated (from external roots) for $\theta>1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho/\theta))$, or $X(S(c,\rho))=X(S(c,\theta\rho))$, respectively.
@@
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. $r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
  • Generalize Def. 2.1 of $\sigma$-soft e/i tests as follows: for $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (so that e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^11] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf. Fig. 1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds. #$\#(D(0,1))<\ell$. $\ell$-testc,ρ, aka $\ell$-test for the disc $D(c,\rho)$, is an $\ell$-test applied to the polynomial $t(y)$ of (1.3).
Observation 6.1.
@@ 7.2 Cauchy sums in the unit disc as weighted power sums of the roots
Cor. 7.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
@@
We can apply it because $|y|=|z|$, while $|z|<1$ by assumption.
@@
Sum equations (7.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 3.1 and (7.2), and obtain
@@
For the roots $z_{j}$ and all $h$, the Cauchy sums $s_{h,q}$ of (7.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$